# TUS-REC2024: A Challenge to Reconstruct 3D Freehand Ultrasound Without External Tracker

### v1

@@ 3.2 Calibration
The calibration process in freehand ultrasound reconstruction involves both spatial and temporal components. Temporal calibration ensures synchronisation between the timestamped ultrasound image frames acquired from the ultrasound machine and the corresponding transformation data recorded by the optical tracking system. This calibration can be performed using the PLUS Toolkit [30], as well as other established methods described in the literature [29, 32, 33, 34]. Synchronisation can be achieved by identifying the optimal time offset that maximises correlation between the probe’s motion (from tracking) and the observed motion in the image stream. Spatial calibration is required to determine the transformation between the ultrasound image coordinate system (in pixels) and the tracker tool coordinate system. Since the optical tracker records the pose of the tracker tool relative to the camera coordinate system, the calibration matrix enables accurate mapping of the ultrasound frames into the spatial coordinate system of the tracking environment.
In this study, a pinhead based method was employed for spatial calibration (as shown in Fig. 1(b)). A pinhead served as the calibration phantom and was repeatedly imaged in the ultrasound images while simultaneously recording the corresponding transformation matrices from the optical tracker. During data acquisition, the ultrasound probe, equipped with tracking markers, was moved at various angles and distances relative to the pinhead. Additionally, the pinhead was designed to appear at different locations within the ultrasound image plane. The calibration was performed in a water medium to ensure optimal ultrasound imaging quality.
The goal of the calibration process is to estimate the transformation matrix that transforms 2D points from the image coordinate system to the camera coordinate system. Let $\{p_{(u_{i},v_{i})}|i=1,...,n\}$ denote the set of 2D coordinates of the pinhead in the ultrasound image coordinate system, and $P$ the corresponding 3D coordinates in the camera coordinate system. Since the pinhead remains stationary throughout the acquisition, each 2D image point, when transformed into 3D space using the estimated calibration matrix, should be expected to converge at the same 3D location $P$, as shown in Eq. (3.2).
$P$ $=$ $T_{1}^{camera\leftarrow tool}\cdot T_{rotation}\cdot T_{scale}\cdot p_{(u_{1},v_{1})}$
$P$ $=$ $T_{2}^{camera\leftarrow tool}\cdot T_{rotation}\cdot T_{scale}\cdot p_{(u_{2},v_{2})}$
@@
$P$ $=$ $T_{n}^{camera\leftarrow tool}\cdot T_{rotation}\cdot T_{scale}\cdot p_{(u_{n},v_{n})}$
where $\{T_{i}^{camera\leftarrow tool}|i=1,...,n\}$ denote the corresponding transformation matrix for each 2D image location of the pinhead, from tracker tool coordinate system to camera coordinate system, recorded from optical camera. The complete calibration matrix is expressed as $T_{calib}=T_{rotation}\cdot T_{scale}$. This composition ensures that pixels are first scaled into physical space, and then mapped into the tracker tool coordinate system using a rigid transformation.
Specifically, in each ultrasound image plane, the 2D location of the pinhead $p_{(u_{i},v_{i})}$ is manually identified in the image coordinate system, while its corresponding physical position in the camera coordinate system, denoted as $P$, remains constant but unknown throughout the acquisition. Consequently, in Eq. (3.2), the parameters to be estimated include the scaling factors ($s_{x}$ and $s_{y}$), the 6-DoF (three rotation angles and three translation components) comprising the rigid transformation $T_{rotation}$, and the 3D position of $P$. These parameters are jointly estimated using a nonlinear least-squares optimisation algorithm [35], in which the objective is to minimise the distance between the transformed 3D locations and the estimated fixed 3D location of the pinhead. The optimisation formulation is given as:
$\mathcal{L}$ $=$ $\min_{T_{rotation},\,T_{scale},\,P}\,\sum_{i=1}^{n}\,dist\,(T_{i}^{camera\leftarrow tool}\cdot T_{rotation}\cdot T_{scale}\cdot p_{(u_{i},v_{i})},P)$ (20)
@@
where $dist\,(\cdot)$ denotes the Euclidean distance computed between corresponding pairs of transformed 3D coordinates and the fixed 3D location $P$.
@@ 3.3 Transformation Estimation
The learning-based freehand ultrasound reconstruction task can be formulated as a pose regression problem, where the goal is to estimate transformations directly from ultrasound frames. The network architecture can vary in terms of its input configuration (e.g., frame pairs or sequences), supervisory label / output representation (rigid or non-rigid; 6-DoF, $4\times 4$ transformation matrix, or points coordinates), and loss function design (in Euclidean space or parameter space).
@@
Input configuration. Given an ultrasound scan, the network can process a sequence of ultrasound frames to estimate the relative transformations between them. A common approach is to use two adjacent frames as input, which can be seen as a specific case of the general sequence input, and predict the relative transformation between them. The 3D ultrasound volume can then be reconstructed by accumulating these estimated relative transformations, as described in Eq. (4). Alternatively, the network can estimate the transformation between non-adjacent frames. This approach enables the network to leverage long-term spatial dependencies, potentially improving robustness and reducing cumulative error in trajectory estimation. A limitation of this approach is that it may not provide transformation estimates for all frames in the sequence.
@@
Supervisory label / Output representation. Since the transformation from the tracker tool to the camera coordinate system, $T_{i}^{camera\leftarrow tool}$, is defined relative to the camera pose, it depends on the external configuration of the camera. This means that scanning the same object from different camera poses produces different transformations (from the tracker tool to the camera coordinate system). As a result, the same input content can be associated with different supervisory labels, which introduces ambiguity and hinders model generalisation. Therefore, transformations that are invariant to the camera’s pose are generally preferred in this application. The supervisory label can be designed as the transformation between two tracker tool coordinate systems corresponding to two ultrasound frames, denoted as $T_{j\leftarrow i}^{tool}$. Alternatively, the calibration matrix can be incorporated into labels to express the transformation in physical space (e.g., in millimeters), $T_{j\leftarrow i}$, or pixel space, via the scaled transformation $T_{scale}^{-1}\cdot T_{j\leftarrow i}\cdot T_{scale}$. It is important to note that all the three supervisory labels, $T_{j\leftarrow i}^{tool}$, $T_{j\leftarrow i}$, and $T_{scale}^{-1}\cdot T_{j\leftarrow i}\cdot T_{scale}$ are rigid transformations. It is crucial to highlight the relationship between the supervisory labels and evaluation metrics discussed in Section 4.3.1, where the supervisory labels serve as ground truth for model training, and the evaluation metrics are employed to assess model performance.
Other examples of the supervisory labels are 6-DoF vector, consisting of three rotation and three translation parameters, and seven-parameter representation using quaternion, comprising four rotation parameters and three translations. These representations are typically derived from the $4\times 4$ rigid transformation matrices, as mentioned above. Additionally, point coordinates derived from Eq. (3) can also be used as supervisory labels. Specifically, the baseline method of this Challenge used the 6-DoF vectors for supervision. It is worth noting that the representation of the network output can be similar to, but does not need to match, the format of the labels used during training. For example, the network may be designed to output a 6-DoF vector, even if the ground truth labels are provided in the form of transformation matrices.
@@
Loss functions. Examples of the loss function include the Euclidean distance between predicted and ground truth point coordinates, as well as the difference between parameters or transformation matrices, regardless of the specific types of labels or predictions. For example, if the ground truth is provided as a $4\times 4$ transformation matrix and the network predicts a 6-DoF vector, the loss can be computed based on the difference between the predicted and ground truth 6-DoF parameters, where the latter are derived from the transformation matrix. Alternatively, the loss can be defined as the point-wise Euclidean distance between points transformed by the ground truth matrix and those transformed by the prediction.

### v2

@@ 2.2 Calibration
The calibration process in freehand ultrasound reconstruction involves both spatial and temporal components. Temporal calibration synchronises timestamps from the ultrasound machine and the optical tracking system, ensuring that the timestamps of each frame are aligned with those of pose measurements. This calibration can be performed using the PLUS Toolkit [19], as well as other established methods described in the literature [18, 21, 22, 23]. For example, synchronisation can be achieved by identifying the optimal time offset that maximises the correlation between the probe’s motion measured by the tracking system and positions of the bottom of a water bath observed in the image stream. Spatial calibration is required to determine the transformation between the ultrasound image coordinate system (in pixels) and the tracker tool coordinate system.
In this study, a pinhead based method was employed for spatial calibration (as shown in Fig. 1(b)). The calibration was performed in a water medium to ensure optimal ultrasound imaging quality. A pinhead sitting under water served as the calibration phantom and was repeatedly imaged in the ultrasound images while the optical tracker simultaneously recorded the corresponding transformation matrices. During data acquisition, the ultrasound probe, equipped with tracking markers, was moved at various angles and distances (typically tens of positions) relative to the pinhead. The pinhead would therefore appear at different locations within the ultrasound image planes.
The goal of the spatial calibration is to estimate the transformation matrix that can transform 2D points in different ultrasound frames to the same point in the fixed camera coordinate system, as well as to estimate the unknown but fixed location of the pinhead. Let $\{p_{(u_{i},v_{i})}|i=1,...,n\}$ denote the set of 2D coordinates of the pinhead in the ultrasound image coordinate system, and $P$ the corresponding 3D coordinates in the camera coordinate system. Since the pinhead remains stationary throughout the acquisition, each 2D image point, when transformed into 3D space using the estimated calibration matrix, should correspond to the same 3D location $P$, as shown in Eq. (2.2).
$P$ $=$ $T_{1}^{camera\leftarrow tool}\cdot T_{rigid}\cdot T_{scale}\cdot p_{(u_{1},v_{1})}$
$P$ $=$ $T_{2}^{camera\leftarrow tool}\cdot T_{rigid}\cdot T_{scale}\cdot p_{(u_{2},v_{2})}$
@@
$P$ $=$ $T_{n}^{camera\leftarrow tool}\cdot T_{rigid}\cdot T_{scale}\cdot p_{(u_{n},v_{n})}$
where $\{T_{i}^{camera\leftarrow tool}|i=1,...,n\}$ denotes the transformation matrices corresponding to each 2D image location of the pinhead, from tracker tool coordinate system to camera coordinate system, recorded from optical tracker. The complete calibration matrix is expressed as $T_{calib}=T_{rigid}\cdot T_{scale}$. This composition ensures that pixels are first scaled into physical unit, and then mapped into the tracker tool coordinate system using a rigid transformation.
Specifically, in each ultrasound frame, the 2D pixel location of the pinhead $p_{(u_{i},v_{i})}$ is manually identified, while its corresponding physical position in the camera coordinate system, denoted as $P$, remains constant but unknown throughout the acquisition. Consequently, in Eq. (2.2), the parameters to be estimated include the scaling factors ($s_{x}$ and $s_{y}$), the 6-DoF (three rotation angles and three translation components) comprising the rigid transformation $T_{rigid}$, and $P$. These parameters are jointly estimated using a nonlinear least-squares optimisation algorithm [24], in which the objective is to minimise the distance between the transformed 3D locations and the estimated fixed 3D location of the pinhead. The optimisation formulation is given as:
$\mathcal{L}$ $=$ $\min_{T_{rigid},\,T_{scale},\,P}\,\sum_{i=1}^{n}\,dist\,(T_{i}^{camera\leftarrow tool}\cdot T_{rigid}\cdot T_{scale}\cdot p_{(u_{i},v_{i})},P)$ (20)
@@
where $dist\,(\cdot)$ denotes the Euclidean distance computed between corresponding pairs of transformed 3D coordinates and $P$.
@@ 2.3 Transformation Estimation
The learning-based trackerless freehand ultrasound reconstruction task can be formulated as a pose regression problem, where the goal is to estimate transformations directly from ultrasound frames. The network architecture can vary in terms of its input configuration (e.g., frame pairs or sequences), output (and its supervisory label) representation (rigid or non-rigid; 6-DoF, $4\times 4$ transformation matrix, or points coordinates), and loss function design (e.g., in Euclidean space or parameter space).
@@
Input configuration. Given an ultrasound scan, the network can process a sequence of ultrasound frames to estimate the relative transformations between them. A common approach is to use two adjacent frames as input, which can be seen as a special case of the general sequence input. The 3D ultrasound volume can then be reconstructed by composing these estimated relative transformations in order, as described in Eq. (4). As the estimation accuracy is influenced by the inter-frame distance, there are also some studies that estimate the transformation between non-adjacent frames. A limitation of this approach is that it may not provide transformation estimations for all frames in the sequence.
@@
Supervisory label / Output representation. Since the transformation from the tracker tool coordinate system to the camera coordinate system, $T_{i}^{camera\leftarrow tool}$, is defined relative to the camera pose, it depends on the external configuration of the camera. This means that scanning the same object from different camera poses produces different transformations (from the tracker tool to the camera coordinate system). As a result, the same set of images can be associated with different supervisory labels, which introduces ambiguity and hinders model generalisation. Therefore, transformations that are invariant to the camera’s pose are generally preferred in this application. The supervisory label can be designed as the transformation between two tracker tool coordinate systems corresponding to two ultrasound frames, denoted as $T_{j\leftarrow i}^{tool}$. Alternatively, the calibration matrix can be incorporated into labels to express the transformation in millimeters, $T_{j\leftarrow i}$, or in pixels, $T_{scale}^{-1}\cdot T_{j\leftarrow i}\cdot T_{scale}$. It is important to note that all the three supervisory labels, $T_{j\leftarrow i}^{tool}$, $T_{j\leftarrow i}$, and $T_{scale}^{-1}\cdot T_{j\leftarrow i}\cdot T_{scale}$, are rigid transformations. It is crucial to highlight the relationship between the supervisory labels and evaluation metrics discussed in Section 4.3.1, where the supervisory labels serve as ground truth for model training, and the evaluation metrics are employed to assess model performance.
Regardless of the form of labels being used, they can be represented as 6-DoF vector, consisting of three rotation and three translation parameters, or 7-DoF vector, comprising four parameters for quaternions and three translations. Additionally, point coordinates derived from Eq. (3) can also be used as supervisory labels.
@@
Loss functions. Depending on the form of supervision label and network output, loss function can vary but in essence should characterise the difference between the output prediction and label. In addition, it is worth noting that the representation of the network output can be similar to, but does not need to match, the format of the labels used during training. For example, if the ground truth is provided as a $4\times 4$ transformation matrix and the network predicts a 6-DoF vector, the loss can be computed based on the difference between the predicted and ground truth 6-DoF parameters, where the latter are derived from the transformation matrix. Alternatively, the loss can be defined as the point-wise Euclidean distance between points transformed by the ground truth matrix and those transformed by the prediction.